# Nested Named Entity Recognition as Single-Pass Sequence Labeling

### v1

@@ Abstract
Alberto Muñoz-Ortiz,1 David Vilares,1 Caio Corro,2 Carlos Gómez-Rodríguez1 1 Universidade da Coruña, CITIC, Spain 2 INSA Rennes, IRISA, Inria, CNRS, Université de Rennes, France {alberto.munoz.ortiz, david.vilares, carlos.gomez}@udc.es caio.corro@insa-rennes.fr
@@ 1 Introduction
Named Entity Recognition (NER) involves identifying token spans that refer to specific named entities. Traditional approaches use the BIO scheme Ramshaw and Marcus (1995), which assigns labels to define the beginning (B), inside (I), and outside (O) tokens of each entity. BIO NER - and other BIO tasks - can be naturally addressed using sequence labeling approaches Ratnaparkhi (1996); Lafferty et al. (2001); Lample et al. (2016); Yang and Zhang (2018). However, nested NER (NNER) challenges these approaches, as it involves entities embedded within other entities, as long as their spans do not overlap. For example, in Figure 1, the person mention “president of the US” contains the geopolitical named entity “the US”. This nested structure breaks the BIO tagging assumption that each token belongs to only one entity. As a result, structured prediction is a better fit for NNER, as it can model interdependent spans and capture hierarchical relationships between nested entities.
@@
Alternatively, some methods transform NNER into constituency parsing by mapping samples to constituent trees and relying on parsing algorithms Finkel and Manning (2009); Wang et al. (2018); Fu et al. (2021); Lou et al. (2022). More recently, Yang and Tu (2022) applied a pointer-network-based bottom-up constituency parser to NNER, offering linear-time decoding. In this context, constituency tree linearization—first introduced by Gómez-Rodríguez and Vilares (2018) and later explored in related work Amini and Cotterell (2022); Kitaev and Klein (2020)—reframes the structured prediction task as a sequence labeling problem. This approach achieves competitive performance while improving computational efficiency, as it outputs the tree in exactly tagging actions. This is in contrast with previous sequence-to-sequence models for NNER where the number of tagging actions depends on the number of predicted mentions Miculicich and Henderson (2020).
Building on this line of work, we present a simple yet effective NNER approach that applies constituency tree linearizations within a standard sequence labeling framework. Our method is single-pass, parser-free, and integrates seamlessly with pretrained encoders. It captures nested structures efficiently and achieves competitive benchmark results, all while requiring minimal implementation effort. Code will be released upon acceptance.
@@ 2 NNER Through Constituent Parsing as Sequence Labeling
We represent NNER structures using constituent trees and adopt a sequence labeling strategy originally developed for syntactic parsing. Our hypothesis is that this formulation aligns well with NNER due to two key properties: (1) the trees are relatively shallow, involving fewer output levels than in full parsing,111In full constituent parsing, large gaps between nonterminals have been shown to challenge sequence tagging approaches Vilares et al. (2019). which eases learning; and (2) the encodings provide full coverage of well-formed trees, regardless of nesting depth. Furthermore, unlike TreeCRF models Fu et al. (2021), span-based Lou et al. (2022) or transition-based parsers Gao et al. (2023), our method avoids complex decoding, enabling linear-time inference— tagging actions—with standard architectures.
@@ 2.1 Linearizations
Let be an input word sequence, where represents the word at position . We define a label for positions , where: (1) is an integer indicating the number of common ancestors between words and , (2) is the non-terminal symbol at their lowest common ancestor, and (3) denotes the unary branch for , if one exists.
In the absolute encoding (abs), represents the number of common ancestors between and . While abs often suffers from a large and sparse label space in full constituent parsing due to the depth and complexity of syntactic trees, we expect it to remain compact and easier to learn in NNER, where structures are shallower.
In the relative encoding (rel), in order to reduce the label set, represents the difference in the number of common ancestors .
Combining these two strategies, the dynamic encoding (dyn) selects, at each position , the most suitable encoding—absolute or relative—for representing the relation between tokens and . It defaults to the relative encoding and switches to absolute when two conditions are met: (1) the relative value , indicating a sharp drop in tree depth with empty non-terminal levels that must be filled in subsequent steps; and (2) the corresponding absolute value , ensuring the constituent node is close to the topmost tree level and the absolute label remains compact, avoiding sparsity and improving learnability. The dynamic encoding uses the absolute function only after abrupt structural changes. This affects few labels—0.75% in ACE2004, 0.73% in ACE2005, 0.15% in GENIA, and 1.16% in NNE—but many sentences: 14.64%, 12.45%, 3.86%, and 20.20%, respectively.
@@ 3 Experiments
We use the MaChAmp framework van der Goot et al. (2021) to train a multitask transformer-based model.222Although the transformer architecture has quadratic complexity, we follow common practice and refer to the rest of the system’s complexity. Specifically, we use roberta-large Liu et al. (2019) and bert-large-uncased Devlin et al. (2019) as shared encoders for the ACE and NNE datasets, and biobert-large-cased-v1.1 Lee et al. (2020) for GENIA. We include RoBERTa as a high-performing option, and BERT for comparison, since it is the most commonly used model in related work. We add three task-specific linear classification heads to predict each atomic component of the label . We adopt the default hyperparameters provided in the MaChAmp repository, as our primary goal is to show these models perform well in off-the-shelf sequence labeling setups with simple plug-and-play scripts. In particular, for encoding and decoding, we use the scripts provided by the CoDeLin library (Roca et al., 2023). This library provides homogeneous post-processing to produce trees from ill-formed label sequences; more details about the post-processing are in Appendix D.
@@
Datasets
Metrics
We evaluate our models in terms of precision, recall, and F1 scores based on strict entity boundaries, requiring both correct spans and types.
3.1 Results
Table 2 reports precision, recall, and F1 scores for our sequence labeling models compared to previous work on NNER. All results are averaged over five runs to provide robust performance estimates (standard deviations are also reported).
Our approach outperforms both comparable and more complex methods on the ACE and NNE datasets. However, this is not the case for GENIA, where the results—particularly recall—are lower. Although the dyn encoding predominantly uses the relative function—as most entities are flat or shallow—it outperforms the other encodings on every dataset except GENIA, effectively leveraging the best of abs and rel linearizations. Although the absolute function is rarely used in the dynamic encoding, it yields small but consistent F1 improvements of 0.74% for ACE2004, 0.39% for ACE2005, 0.22% for GENIA, and 0.02% for NNE. As more than 99% of entities in GENIA are flat or singly nested, the dyn and rel encodings represent the entities almost identically. Tetra-tagging is the worst-performing encoding in all but one setup.
Results by depth
Table 3 presents precision and recall by depth for each encoding and dataset. Since the number of predicted entities may vary across runs, we follow Shibuya and Hovy (2020) and report the precision results from the run whose F1 is closest to the average, and the average recall over five runs. abs performs best for flat entities on all datasets, dyn leads for singly nested entities in ACE2004 and ACE2005, rel in GENIA, and 4tg and dyn tie in NNE. For entities nested two or more times, abs performs best in GENIA and achieves the highest precision on the ACE datasets, but is outperformed in recall except in GENIA. In NNE, 4tg achieves the highest recall.
@@ 3.1 Results
Results by entity length
Table 2 presents precision and recall by span length. Among the encodings, rel and dyn perform best on single-token spans and those of 2–4 tokens. Results for 5–9 token spans are mixed, while abs consistently outperforms the others on longer spans (10 tokens), except in ACE2005, where it is surpassed by dyn. 4tg performs well on short spans—particularly in NNE—but its performance drops sharply as span length increases. These trends suggest that dyn is well-suited for frequent, shorter entities, while abs offers greater robustness on longer ones.
@@ 4 Conclusion
We cast nested NER for the first time as a single-pass sequence labeling task by first transforming NNER annotations into constituent trees, and then leveraging the linearization options these trees enable. In particular, we explored both depth-based (absolute, relative and dynamic encodings) and transition-based encodings (tetra-tagging) previously proposed for this type of tree structure. Our experiments show that this approach models nested structures effectively, with competitive performance and no need for complex architectures—remaining lightweight and compatible with standard sequence labeling tools.
Among the explored encodings, the dynamic variant—which selects between absolute and relative schemes based on local structural cues—consistently achieves the highest F1 across datasets, except for GENIA. The transition-based tetra-tagging approach remains competitive on flat and short entities but degrades on longer spans, where depth-based strategies prove more effective.
@@ Acknowledgements
This work was funded by SCANNER-UDC (PID2020-113230RB-C21) funded by MICIU/AEI/10.13039/501100011033; Xunta de Galicia (ED431C 2024/02); GAP (PID2022-139308OA-I00) funded by MICIU/AEI/10.13039/501100011033/ and by ERDF, EU; Grant PRE2021-097001 funded by MICIU/AEI/10.13039/501100011033 and by ESF+ (predoctoral training grant associated to project PID2020-113230RB-C21); LATCHING (PID2023-147129OB-C21) funded by MICIU/AEI/10.13039/501100011033 and ERDF; and Centro de Investigación de Galicia “CITIC”, funded by the Xunta de Galicia through the collaboration agreement between the Consellería de Cultura, Educación, Formación Profesional e Universidades and the Galician universities for the reinforcement of the research centres of the Galician University System (CIGUS).
@@ Appendix A Precision and Recall Results
We now show in Table 5 the precision and recall corresponding to the evaluation in Table 2 of the main text.
@@
Appendix B Dataset information
Table 6 shows statistics about each dataset, including data on the depth and length of entities.
@@ Appendix B Dataset information
Appendix C Data coverage
The absolute, relative, and dynamic encodings are able to represent all nested entity structures in the datasets perfectly. The only unrecoverable cases under these encodings stem from annotation inconsistencies—such as crossing spans in ACE2004 or inverted boundaries in NNE—and not from encoding limitations. The tetra-tagging encoding (4tg), while equally expressive in theory, exhibits a small number of unrecoverable entities due to decoding ambiguities in the CoDeLin implementation rather than the encoding design itself. Table 7 reports the maximum recall for each encoding and dataset, with the number of unrecoverable entities shown in parentheses.
| ACE2004 | ACE2005 | GENIA | NNE | |
| abs | 99.97 (8) | 100.00 | 100.00 | >99.99 (1) |
| rel | 99.97 (8) | 100.00 | 100.00 | >99.99 (1) |
| dyn | 99.97 (8) | 100.00 | 100.00 | >99.99 (1) |
| 4tg | 99.97 (9) | 99.99 (2) | 100.00 | 99.96 (118) |
Appendix D Post-processing
Rather than imposing strict constraints during prediction, parsing as sequence labeling ensures well-formedness by applying simple heuristics to ill-formed label sequences. In our setup we apply the heuristics provided by CoDeLin (Roca et al., 2023), i.e., (1) in depth-based encodings, if multiple labels are assigned to the same non-terminal node, we retain only the first; (2) if no label is assigned to a node, it is removed from the tree; (3) indexes out of range (e.g. -7 in the relative encoding when there are fewer than 7 levels to go up) are changed to the nearest legal index; and (4) in tetra-tagging, if the label sequence specifies an invalid transition, the transition is skipped. Words that remain unattached due to skipped transitions are attached to the lowest nonterminal in the rightmost tree spine.
Appendix E Model Size and Budget
We fine-tune three models: BERT Large (340M parameters), BioBERT Large (340M), and RoBERTa Large (355M). Each model is trained with 4 encoding strategies, across 4 datasets, and using 5 random seeds, resulting in a total of 240 training runs. Training was performed on a GPU cluster using NVIDIA A100 GPUs (40GB), with each run executed on a single GPU. The cumulative training time is estimated at approximately 300–360 GPU hours.

### v2

@@ Abstract
Alberto Muñoz‐Ortiz,1 David Vilares,1 Caio Corro,2 Carlos Gómez‐Rodríguez1 1 Universidade da Coruña, CITIC, Spain 2 INSA Rennes, IRISA, Inria, CNRS, Université de Rennes, France {alberto.munoz.ortiz, david.vilares, carlos.gomez}@udc.es caio.corro@insa‐rennes.fr
@@ 1 Introduction
Named Entity Recognition (NER) involves identifying token spans that refer to specific named entities. Traditional approaches use the BIO scheme Ramshaw and Marcus (1995), which assigns labels to define the beginning (B), inside (I), and outside (O) tokens of each entity. BIO NER—and other BIO tasks—can be naturally addressed using sequence labeling approaches Ratnaparkhi (1996); Lafferty et al. (2001); Lample et al. (2016); Yang and Zhang (2018). However, nested NER (NNER) challenges these approaches, as it involves entities embedded within other entities as long as their spans do not overlap. For example, in Figure 1, the person mention “president of the USA” contains the geopolitical named entity “the USA”. This nested structure breaks the BIO tagging assumption that each token belongs to only one entity. As a result, structured prediction is a better fit for NNER, as it can model interdependent spans and capture hierarchical relationships between nested entities.
@@
Alternatively, some methods transform NNER into constituency parsing by mapping samples to constituent trees and relying on parsing algorithms Finkel and Manning (2009); Wang et al. (2018); Fu et al. (2021); Lou et al. (2022). More recently, Yang and Tu (2022) applied a pointer-network-based bottom-up constituency parser to NNER, offering linear-time decoding. In this context, constituency tree linearization—first introduced by Gómez-Rodríguez and Vilares (2018) and later explored in related work Kitaev and Klein (2020); Amini and Cotterell (2022)—reframes the structured prediction task as a sequence labeling problem. This approach achieves competitive performance while improving computational efficiency, as it outputs the tree in exactly tagging actions. This is in contrast with previous sequence-to-sequence models for NNER where the number of tagging actions depends on the number of predicted mentions Miculicich and Henderson (2020).
Building on this line of work, we present a simple yet effective NNER approach that applies constituency tree linearizations within a standard sequence labeling framework.111Code is available at https://github.com/amunozo/nner_as_sl. Our method is single-pass, parser-free, and integrates seamlessly with pretrained encoders. It captures nested structures efficiently and achieves competitive benchmark results, all while requiring minimal implementation effort. Overall, the data preparation remains lightweight and fully automated, and it is no more complex than in traditional nested NER settings, thus preserving the simplicity of standard NER pipelines while adding structural expressiveness.
@@ 2 NNER Through Constituent Parsing as Sequence Labeling
We represent NNER structures using constituent trees, adopting a sequence labeling strategy for syntactic parsing. We hypothesize that this formulation aligns well with NNER due to two key properties: (1) the trees are relatively shallow, involving fewer output levels than in full parsing,222In full constituent parsing, large gaps between nonterminals have been shown to challenge sequence tagging approaches Vilares et al. (2019). and (2) the encodings provide full coverage of well-formed trees, regardless of nesting depth. Furthermore, unlike TreeCRF models Fu et al. (2021), span-based Lou et al. (2022) or transition-based parsers Gao et al. (2023), our method avoids complex decoding, enabling linear-time inference333Although the transformer architecture has quadratic complexity, we follow common practice and refer to the rest of the system’s complexity Shibuya and Hovy (2020); Corro (2023); Lou et al. (2022).— tagging actions—with standard architectures.
@@ 2.1 Linearizations
Let be an input word sequence, where represents the word at position . We define a label for each word , where: (1) is an integer indicating the number of common ancestors between words and , (2) is the non-terminal symbol at their lowest common ancestor, and (3) denotes the unary branch for , if one exists.
In the absolute encoding (abs) (Gómez-Rodríguez and Vilares, 2018), represents the number of common ancestors between and . While abs often suffers from a large and sparse label space in full constituent parsing due to the depth and complexity of syntactic trees, we expect it to remain compact and easier to learn in NNER, where structures are shallower.
The relative encoding (rel) was also introduced in (Gómez-Rodríguez and Vilares, 2018). In order to reduce the label set, represents the difference in the number of common ancestors, i.e., for (and ).
Combining these two strategies, the dynamic encoding (dyn) (Vilares et al., 2019) selects, at each position , the most suitable encoding—absolute or relative—for representing the relation between tokens and . It defaults to the relative encoding and switches to absolute when two conditions are met: (1) the relative value , indicating a sharp drop in tree depth with empty non-terminal levels that must be filled in subsequent steps; and (2) the corresponding absolute value , ensuring the constituent node is close to the topmost tree level and the absolute label remains compact, avoiding sparsity and improving learnability. The dynamic encoding uses the absolute function only after abrupt structural changes. This affects few labels—0.75% in ACE2004, 0.73% in ACE2005, 0.15% in GENIA, and 1.16% in NNE—but many sentences: 14.64%, 12.45%, 3.86%, and 20.20%, respectively.
@@ 3 Experiments
We use the MaChAmp framework van der Goot et al. (2021) to train a multitask transformer-based model. Specifically, we use roberta-large Liu et al. (2019) and bert-large-uncased Devlin et al. (2019) as shared encoders for the ACE and NNE datasets, and biobert-large-cased-v1.1 Lee et al. (2020) for GENIA. We include RoBERTa as a high-performing option, and BERT for comparison, since it is the most commonly used model in related work. We add three task-specific linear classification heads to predict each atomic component of the label . We adopt the default hyperparameters provided in the MaChAmp repository, as our primary goal is to show that these models perform well in off-the-shelf sequence labeling setups with simple plug-and-play scripts. In particular, for encoding and decoding, we use the scripts provided by the CoDeLin library (Roca et al., 2023). This library provides homogeneous post-processing to produce trees from ill-formed label sequences; more details about the post-processing are in Appendix E.
@@
| S20 | 0 | – | – | 80.4 | 76.1 | 80.3 | 77.9 | – | – |
| 1 | – | – | 72.4 | 71.7 | 57.1 | 40.6 | – | – | |
| 1 | – | – | 73.2 | 70.4 | 57.1 | 40.6 | – | – | |
| 2 | – | – | 77.6 | 57.5 | 66.7 | 0.0 | – | – | |
Datasets
Metrics
We evaluate our models in terms of precision, recall, and F1-scores based on strict entity boundaries, requiring both correct spans and types. We also provide finer-grained evaluations (e.g., by entity length) and external comparisons.
3.1 Results
Table 2 reports precision, recall, and F1-scores for our sequence labeling models compared to previous work on NNER. All results are averaged over five runs to provide robust performance estimates (standard deviations are also reported).
Our approach outperforms both comparable and more complex methods on the ACE and NNE datasets. However, this is not the case for GENIA, where the results—particularly recall—are lower. Although the dyn encoding predominantly uses the relative function—as most entities are flat or shallow—it outperforms the other encodings on every dataset except GENIA, effectively leveraging the best of abs and rel linearizations. Even though the absolute function is rarely used in the dynamic encoding, it yields small but consistent F1 improvements of 0.74% for ACE2004, 0.39% for ACE2005, 0.22% for GENIA, and 0.02% for NNE. As more than 99% of entities in GENIA are flat or singly nested, the dyn and rel encodings represent the entities almost identically. Tetra-tagging is the worst-performing encoding in all but one setup.
@@ 3.1 Results
| S21 | 1 | 89.6 | 91.0 | – | – | – | – | – | – |
| 2 | 85.3 | 83.7 | – | – | – | – | – | – | |
| 2–4 | 87.0 | 86.0 | – | – | – | – | – | – | |
| 5–9 | 84.3 | 83.2 | – | – | – | – | – | – | |
| 10 | 69.2 | 63.2 | – | – | – | – | – | – | |
| ACE2004 | ACE2005 | GENIA | NNE | |||||||||||||
| Best | Worst | Mean | Std | Best | Worst | Mean | Std | Best | Worst | Mean | Std | Best | Worst | Mean | Std | |
| 4tg | PER | FAC | 78.9 | 10.6 | PER | LOC | 77.8 | 7.4 | protein | DNA | 73.1 | 3.0 | ENERGY | BAND | 73.7 | 29.8 |
| abs | PER | WEA | 78.5 | 10.8 | PER | LOC | 78.6 | 7.2 | protein | cell_type | 74.7 | 3.4 | DISEASE | BAND | 74.4 | 29.5 |
| dyn | PER | FAC | 81.3 | 8.3 | PER | LOC | 79.6 | 7.6 | protein | DNA | 73.1 | 3.2 | ENERGY | BAND | 74.0 | 29.4 |
| rel | PER | FAC | 80.6 | 8.8 | PER | LOC | 79.3 | 7.7 | protein | DNA | 73.0 | 3.2 | HURRICANE | BAND | 73.4 | 30.4 |
Results by depth
Table 3 presents precision and recall by depth for each encoding and dataset. Since the number of predicted entities may vary across runs, we follow Shibuya and Hovy (2020) and report the precision results from the run whose F1 is closest to the average, and the average recall over five runs. abs performs best for flat entities on all datasets, dyn leads for singly nested entities in ACE2004 and ACE2005, rel in GENIA, and 4tg and dyn tie in NNE. For entities nested two or more times, abs performs best in GENIA and achieves the highest precision on the ACE datasets, but is outperformed in recall except in GENIA. In NNE, 4tg achieves the highest recall. We also compare against external work that has reported these results, particularly Shibuya and Hovy (2020), where our method shows superior performance on flat entities but underperforms on deep nested structures.
Results by entity length
Table 4 presents precision and recall by span length. Among the encodings, rel and dyn tend to perform best on single-token spans and those of 2–4 tokens. Results for 5–9 token spans are mixed, while abs consistently outperforms the others on longer spans (10 tokens), except in ACE2005, where it is surpassed by dyn. 4tg performs well on short spans—particularly in NNE—but its performance drops sharply as span length increases. These trends suggest that dyn is well-suited for frequent, shorter entities, while abs offers greater robustness on longer ones. In addition, we benchmark against external work we are aware of, such as the more complex model from Shen et al. (2021), where our methods are competitive for entities spanning fewer than four tokens, but the performance gap widens for longer entities where their specialized architecture appears more advantageous.
Results by entity type
Table 5 provides a detailed analysis of F1-scores for each entity type, identifying the best and worst performers and summarizing the overall distribution with mean and standard deviation. PER (Person) consistently achieves the highest scores in both ACE datasets, while protein is the strongest performer in GENIA, regardless of the encoding used.
Focusing on the variance, scores in GENIA and ACE2005 are stable across types, with a standard deviation of around 3 and 7.5 respectively for all encodings. For ACE2004, the rel and dyn encodings show a standard deviation of approximately 8.5, while 4tg and abs present slightly higher variation at around 10. As expected, NNE shows a much higher standard deviation (29.7), due to its large and sparse label set; however, this pattern is stable across encodings.
3.2 Discussion
The results show that with standard transformer-based encoders, NNER can be effectively learned as sequence labeling under various representations. As expected, for both flat and short entities, differences across encodings do not appear to substantially affect performance, since in these cases representing sentences as constituent trees does not lead to significant changes. For deep entities, more elaborate encodings appear to be beneficial compared to the naive absolute encoding. For longer entities, results are mixed. This is likely influenced by the fact that longer entities can be either flat or deep, which may confound conclusions about the impact of different encodings. In particular, depth-based encodings perform better on the ACE and GENIA datasets, while 4tg performs comparably to the other encodings on the NNE dataset. In this context, a factor could be the relative ease of the NNE dataset. Although it includes many nested entities and a large label set, it also provides a substantial larger amount of training data (see Table 8). If the learning task becomes less demanding overall, differences between encodings are less pronounced, and the encoder may play a more prominent role. As shown by our NNE results, the narrow F1 range (94.0–94.6) across encodings and models suggests that, with enough data, different encodings are learned to similar capacity, thereby limiting observable performance differences.
@@ 4 Conclusion
We cast nested NER for the first time as a single-pass sequence labeling task by first transforming NNER annotations into constituent trees, and then leveraging the linearization options these trees enable. In particular, we explored both depth-based (absolute, relative and dynamic encodings) and transition-based encodings (tetra-tagging) previously proposed for this type of tree structure. Our experiments show that this approach models nested structures effectively, with competitive performance and no need for complex architectures, remaining lightweight and compatible with standard sequence labeling tools.
Among the explored encodings, the dynamic variant—which selects between absolute and relative schemes based on local structural cues—consistently achieves the highest F1 across datasets, except for GENIA. The transition-based tetra-tagging approach is competitive on flat and short entities but degrades on longer spans, where depth-based strategies prove more effective.
@@ Acknowledgements
This work was funded by SCANNER-UDC (PID2020-113230RB-C21) funded by MICIU/AEI/10.13039/501100011033; Xunta de Galicia (ED431C 2024/02); GAP (PID2022-139308OA-I00) funded by MICIU/AEI/10.13039/501100011033/ and by ERDF, EU; Grant PRE2021-097001 funded by MICIU/AEI/10.13039/501100011033 and by ESF+ (predoctoral training grant associated to project PID2020-113230RB-C21); LATCHING (PID2023-147129OB-C21) funded by MICIU/AEI/10.13039/501100011033 and ERDF; TSI-100925-2023-1 funded by Ministry for Digital Transformation and Civil Service and “NextGenerationEU” PRTR; and Centro de Investigación de Galicia “CITIC”, funded by the Xunta de Galicia through the collaboration agreement between the Consellería de Cultura, Educación, Formación Profesional e Universidades and the Galician universities for the reinforcement of the research centres of the Galician University System (CIGUSA).
@@ Appendix A Precision and Recall Results
Table 6 shows the precision and recall corresponding to the evaluation presented in Table 2 of the main text.
@@
Appendix B Results from Best Runs
Table 7 presents the precision, recall, and F1-scores from the best run of each model, selected according to the highest F1-score.
| ACE2004 | ACE2005 | GENIA | NNE | |||||||||
| Model | Prec. | Rec. | F1 | Prec. | Rec. | F1 | Prec. | Rec. | F1 | Prec. | Rec. | F1 |
| abs | 86.4 | 86.6 | 86.5 | 83.7 | 85.7 | 84.7 | 77.1 | 75.5 | 76.3 | 93.6 | 93.8 | 93.7 |
| rel | 86.9 | 86.1 | 86.5 | 85.1 | 85.1 | 85.1 | 79.3 | 69.8 | 74.3 | 94.3 | 93.9 | 94.1 |
| dyn | 87.7 | 87.2 | 87.4 | 85.2 | 85.7 | 85.5 | 78.5 | 71.4 | 74.8 | 93.9 | 93.9 | 93.9 |
| 4tg | 83.8 | 86.6 | 85.1 | 83.1 | 84.7 | 83.9 | 74.6 | 72.6 | 73.5 | 93.7 | 94.2 | 93.9 |
Appendix C Dataset information
Table 8 presents statistics for each dataset, including information on entity depth and length.
@@ Appendix C Dataset information
Appendix D Data coverage
The encodings can represent all nested entity structures in the datasets without error. The only unrecoverable cases with these encodings arise from annotation inconsistencies—such as crossing spans in ACE2004 or inverted boundaries in NNE—rather than from limitations of the encodings themselves.
Appendix E Post-processing
Rather than imposing strict constraints during prediction, parsing as sequence labeling ensures well-formedness by applying simple heuristics to ill-formed label sequences. In our setup, we apply the heuristics provided by CoDeLin (Roca et al., 2023), i.e., (1) in depth-based encodings, if multiple labels are assigned to the same non-terminal node, we retain only the first; (2) if no label is assigned to a node, it is removed from the tree; (3) indices out of range (e.g., -7 in the relative encoding when there are fewer than 7 levels to go up) are changed to the nearest legal index; and (4) in tetra-tagging, if the label sequence specifies an invalid transition, the transition is skipped. Words that remain unattached due to skipped transitions are attached to the lowest non-terminal in the rightmost tree spine.
Appendix F Model Size and Budget
We fine-tune three models: BERT Large (340M parameters), BioBERT Large (340M), and RoBERTa Large (355M). Each model is trained with four encoding strategies, across four datasets, and using five random seeds, resulting in a total of 240 training runs. Training was performed on a GPU cluster using NVIDIA A100 GPUs (40GB), with each run executed on a single GPU. The cumulative training time is estimated at approximately 300–360 GPU hours.